Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O3)
menzione - 22:55 Wednesday 10 July 2019 (46321) Print this report
Operator Report - Afternoon shift

ITF found in Troubleshooting mode, It has been recovered at 16:00 UTC (except SQZ) by Casanueva, Gouaty, Vardaro, Was, Tacca.
When ITF locked in Low_Noise3 (without SQZ) at 16:10 UTC the BNS Range was~40 Mpc with big drops.
Upon stibilized (at 17:26 UTC), ITF set again in Science Mode. BNS Range ~43 Mpc. Stable lock.
ITF unlocked due to a fast unlock (plot attached) at 20:44 UTC. Relocking in progress.

GRB
15:38 UTC - [gracedb] Fermi event created: E338278. Flag reset at 15:40 UTC. (ITF down)

Guard Tour
17:32 - 18:05 UTC
19:36 - 20:07 UTC
 

Calibration
According with DET experts and Commissioning Coordinator, the standard Calibrations and Longitudinal Noise Injections have been postponed due to the troubleshooting on DET.

DMS
- FoxBox is temporarily out of order, "ping(foxbox1)" SubFlag shalved till Friday.

ISYS
PSTAB
13:03 - 13:05 UTC: Open and Close the PSTAB in order to remove the extra noise. Action performed also at:
13:43 - 13:44 UTC

SUSP
At 13:10 UTC the WE Guardians opened LC. Properly closed.

DET
(09-07-2019 11:30 - 10-07-2019 17:25) Photodiode failure - ITF in Troubleshooting Mode

Images attached to this report
Comments to this report:
minazzoli - 9:14 Thursday 11 July 2019 (46326) Print this report

The fast unlock indeed seems to be a regular fast unlock (plot 1). With an increase of the number of (essentially negative) glitches in EOM_CORR a second before the unlock (plot2).

Images attached to this comment
verkindt - 14:19 Thursday 11 July 2019 (46327) Print this report

As mentioned by Olivier, just before this fast unlock, the INJ_EOM_CORR_raw_FS signal had some negative value glitches.
Plot 1 shows a zoom of the last three glitches about 1 ms before the unlock.
Plot 2 is a zoom on the unlock, in case it can help investigations.

Images attached to this comment
Search Help
×

Warning

×